# Adaptive Modality Balanced Online Knowledge Distillation for Brain-Eye-Computer based Dim Object Detection

### v2

@@ V Experimental Results
In this section, the experimental settings with the parameter configurations of our proposed approach are first presented. Subsequently, a comprehensive comparison and ablation study are conducted to evaluate the performance of the proposed AMBOKD approach in the dim object recognition task. Then, the effectiveness and generalisability of the OKD and AMB modules incorporated in AMBOKD are evaluated. In addition, the transfer experiments and system verifications by simulating real scenarios are designed to further verify the generalization performance of the proposed method and system under few-shot conditions.
@@ V-A1 Configurations
The visual encoder uses EfficientNet [22] as the basic approach for feature extraction from the input images. Furthermore, for ablation analysis, ResNet-50 [18] and MobileNetV2 [19] are used as alternative visual encoders. The EEG encoder is based on our previous study called MCGRAM [8], which is employed to extract features from the EEG data. In addition, TSception [58] and EEGNet [57] are used as alternative EEG encoders for the ablation analysis. The parameters of the visual and the EEG encoders are set to follow the configurations as described in the previous study, providing consistency and comparability with the existing studies. To align the features obtained from the visual and EEG encoders, FC layers are used to scale the lengths of both sets of features to a common length of 64. This ensures compatibility for subsequent processing and fusion of the multimodal features.
The number of heads is set to 2 in the multi-head self-attention module of the fusion model. The kernel sizes of the FC layers for alignment are configured as $1280\times 64$ and $256\times 64$ for visual and EEG features, respectively. Before computing the Q, K, and V matrices, the matrices in the FC layer have a size of $64\times 64$. The FC layers have a size of $128\times 2$ in the classifier of the Fusion domain.
The temperature coefficient $\tau$ is set to 4 in the OKD approach to control the softness of the KD targets. The hyperparameters $\alpha$ and $\beta$ are dynamically adjusted by the AMB module, which contributes to the regularization component of the overall loss function. The hyperparameter $\gamma$ is set to 3 to fine-tune the sensitivity of the AMB module. In addition, we impose lower and upper limits, $R_{\min}$ and $R_{\max}$ at 0.1 and 10, respectively, to confine the range of dynamic modulation ratio $R^{DG}_{(t)}$ during the training process.
The PyTorch framework is used to implement the AMBOKD model. The Adam optimizer with a learning rate of $10^{-4}$ is used to optimize all network parameters. The model is trained for 15 epochs with a batch size of 64. The optimization loss function employed comprised the CE loss and the KD loss. A cross-validation approach is used to assess the performance of the AMBOKD model. The data of each subject are chosen as the validation set, whereas the data of the remaining nine subjects are employed as the training set. We repeat this process for each fold with five different seeds, and we obtain the final experimental finding by averaging the results of all the folds and seeds.
@@ V-A2 Performance Metrics
A comprehensive set of metrics, including area under the receiver operating characteristic curve (AUC), accuracy (ACC), F1 score (F1) and Precision are used to evaluate the performance of classification models. Each metric is analyzed in terms of mean and standard deviation, offering insight into the reliability of the model and its ability to generalize across various datasets or conditions. Specifically, the AUC is emphasized as the main metric because of its effectiveness in measuring the discriminative ability of a model, especially in scenarios with unbalanced class distributions. To evaluate the object detection performance of the brain-eye-computer based system, AP@50:95 and AR@50:95 are used to measure the average precision and recall over Intersection Over Union (IOU) thresholds ranging from 0.50 to 0.95, while AP@50 and AR@50 are used for an IOU of 0.50. In addition, F1@50:95 and F1@50 provide a balanced view of accuracy by considering both precision and recall at these thresholds.
TABLE II: The subject-dependent results of the baseline and state-of-art methods in comparison studies.
Session Method AUC/Std (%)
Subject 1 Subject 2 Subject 3 Subject 4 Subject 5 Subject 6 Subject 7 Subject 8 Subject 9 Subject 10 Average
1 EEGNet 63.57/6.44 64.12/6.68 61.60/7.72 54.55/7.16 60.14/8.58 61.49/7.17 54.22/5.27 61.66/6.21 69.59/7.25 56.30/6.83 60.72/6.93
MCGRAM 81.67/7.49 87.10/5.54 86.36/6.09 78.20/9.01 80.22/6.75 82.31/5.68 77.47/6.90 79.59/6.15 91.79/4.71 78.29/5.65 82.30/6.40
AMBOKD-E* 86.98/6.35 90.57/4.70 88.78/6.01 81.64/7.71 87.30/4.85 88.33/4.83 81.41/5.78 86.18/4.50 93.66/3.80 82.56/4.63 86.74/5.32
ResNet-50 83.22/6.03 81.73/6.00 81.67/5.53 79.56/6.87 80.75/6.48 77.86/7.12 83.37/6.43 79.29/6.69 81.68/5.75 80.85/6.42 81.00/6.33
EfficientNet 76.28/9.97 74.91/10.87 79.74/8.73 79.27/8.30 76.99/8.40 73.59/10.40 75.54/8.19 76.70/8.96 78.75/8.27 73.09/9.43 76.49/9.15
AMBOKD-V* 78.28/10.18 82.63/8.81 86.31/6.56 82.39/7.88 83.74/5.84 81.64/7.09 80.86/8.42 82.18/6.91 79.97/7.04 78.67/8.89 81.67/7.76
MLB 89.82/4.73 87.44/5.40 92.07/5.05 84.22/8.33 86.28/6.54 75.79/9.03 83.75/6.79 83.18/7.33 92.33/4.46 83.86/7.21 85.87/6.49
MKD 89.31/5.74 89.51/5.12 92.87/4.38 85.81/7.00 88.46/5.32 82.62/6.12 87.41/6.28 87.59/5.12 93.46/3.45 85.32/6.81 88.24/5.53
EMKD 90.38/4.99 89.56/5.15 92.93/4.30 85.73/6.72 88.70/5.15 82.20/6.06 86.62/6.62 86.38/5.89 94.06/3.24 85.28/6.92 88.18/5.50
CA-MKD 90.23/4.97 89.62/5.12 92.45/4.43 85.93/7.29 87.52/5.81 81.72/7.05 86.17/6.43 87.72/4.90 93.11/3.98 85.62/6.31 88.01/5.63
DML* 91.02/4.73 89.68/4.61 92.44/4.35 86.63/7.18 88.35/5.49 89.14/4.49 87.64/5.70 89.45/4.22 94.07/3.28 87.73/5.43 89.62/4.95
KDCL* 91.04/5.00 89.45/4.97 91.51/5.11 86.35/7.20 87.30/6.23 87.82/5.71 86.91/6.12 87.43/4.60 93.83/3.82 85.23/7.05 88.69/5.58
EML* 88.50/5.88 83.49/7.64 87.76/5.82 83.82/7.27 84.06/7.61 85.95/6.25 81.47/8.63 84.48/5.73 90.63/4.44 80.23/7.66 85.04/6.69
AMBOKD (ours)* 94.29/3.89 94.29/3.63 95.37/3.30 89.56/5.53 93.99/3.49 94.50/3.30 92.34/5.25 93.80/3.52 96.69/2.64 91.78/3.56 93.63/3.81
2 EEGNet 58.40/6.78 66.72/6.18 55.52/5.99 59.68/7.43 60.74/8.02 57.46/6.14 58.15/8.59 67.20/6.64 62.72/7.64 57.44/6.60 60.40/7.00
MCGRAM 69.29/8.21 86.73/5.49 72.86/6.83 79.39/7.15 79.54/7.29 78.40/7.25 74.39/7.99 86.32/4.77 90.01/4.01 77.03/5.56 79.40/6.46
AMBOKD-E* 74.87/7.51 90.37/4.73 78.00/6.94 87.15/4.67 85.99/6.60 82.78/7.19 79.37/6.62 90.21/3.56 92.50/3.17 83.13/5.11 88.39/5.61
ResNet-50 83.49/6.72 80.59/6.89 82.76/6.37 79.72/5.42 84.70/5.41 79.40/6.48 80.23/6.51 82.58/6.40 78.43/5.51 80.55/6.37 81.25/6.21
EfficientNet 73.77/9.28 80.58/8.90 73.05/10.96 73.54/11.58 74.76/10.13 75.96/10.85 74.12/8.66 74.42/8.60 75.33/10.74 72.65/10.40 74.82/10.01
AMBOKD-V* 85.02/6.04 86.06/6.74 86.33/72.74 83.79/6.95 82.37/7.33 83.66/7.08 78.47/72.58 74.87/7.78 79.45/7.52 77.37/8.52 81.74/7.25
MLB 78.24/8.62 90.62/6.12 80.23/6.53 85.27/5.69 82.99/7.49 77.27/8.72 85.00/5.92 83.19/5.93 89.99/5.39 85.18/5.83 83.80/6.62
MKD 81.98/7.78 92.82/3.91 79.38/7.58 84.04/6.68 85.08/6.22 81.48/7.88 85.15/6.48 87.82/5.19 91.60/4.86 85.29/6.52 85.46/6.31
EMKD 82.38/8.14 93.69/3.66 80.53/6.96 84.86/6.43 85.53/6.13 78.71/8.58 85.12/6.50 87.12/5.97 91.88/4.50 84.75/6.69 85.46/6.36
CA-MKD 86.01/8.41 94.19/3.82 85.09/5.96 84.16/6.64 84.05/6.99 79.97/8.15 85.35/6.22 87.40/5.16 93.72/3.69 84.40/7.17 86.43/6.22
DML* 81.09/8.21 89.87/4.85 81.62/6.94 85.18/5.94 88.33/5.25 87.97/4.57 85.89/6.29 89.82/3.82 92.58/4.01 83.76/6.28 86.61/5.62
KDCL* 79.15/8.96 86.05/6.04 78.99/6.65 83.08/7.13 86.93/5.31 87.12/4.45 84.96/6.85 88.41/3.94 92.74/4.12 85.20/6.32 85.26/5.98
EML* 81.09/8.18 83.99/7.24 79.19/6.87 81.25/7.49 87.51/4.86 85.89/6.38 83.90/6.09 87.89/4.76 89.21/5.30 83.53/6.52 84.34/6.37
AMBOKD (ours)* 89.56/5.20 96.97/2.40 90.37/4.76 93.94/3.61 94.09/4.00 93.08/3.99 90.54/4.64 95.33/2.44 96.58/2.99 91.83/3.96 93.23/3.80
* Online knowledge distillation method
@@ V-B1 Comparison Results on ESSVP Dataset
To verify the effectiveness of the proposed AMBOKD approach, we conduct a series of comparison experiments with the baseline and state-of-art models, including unimodal approaches EEGNet [57], MCGRAM [8], ResNet-50 [18], EfficientNet [22], multimodal fusion approach MLB [62], multimodal knowledge distillation approaches MKD [49], EMKD [51], CA-MKD [50], multimodal online knowledge distillation approaches DML [47], KDCL [52], EML [53]. In addition, the EEG and visual modal model trained in AMBOKD are named AMBOKD-E and AMBOKD-V.
As indicated in Table I, our proposed AMBOKD achieve impressive results in the dim object recognition task of subject-independent experiments. This outcome surpasses the performance of baseline and state-of-the-art approaches. Meanwhile, the mutual learning mode with AMB module enhances the performance of EEG and visual modality models: all three metrics of AMBOKD-E and AMBOKD-V surpass those of the original unimodality models MCGRAM and Efficient-Net. The AUC-based ANOVA statistical test findings demonstrate that our proposed AMBOKD, AMBOKD-E, and AMBOKD-V significantly outperform all comparison approaches in subject-independent experiments ($p<0.05$)
Table II shows the comparison results obtained from subject-dependent experiments. Notably, optimal results are achieved across all subjects using our approach, indicating significant advantages. This finding serves as a compelling demonstration of the effectiveness and adaptability of our method when applied to subject-specific experiments within practical scenarios.
@@ V-B2 Comparison Results on CIFAR-100 Dataset
We use the same training setup as described in [64], employing an SGD optimizer with a momentum of 0.9, a batchsize of 64, a weight decay of $5\times 10^{-4}$. The initial learning rate is set at 0.05, with a decay factor of 10 every 30 epochs starting from epoch 150. For the AMB module, we specify a learning rate of $1\times 10^{-4}$. The comparison results with three different random seeds, presented as acc(@1, mean/standard deviation), are shown in Table III. Our AMBOKD achieve the best performance among all compared methods, demonstrating its robustness and practicability.
TABLE III: Comparison results on CIFAR-100 Test set
@@ V-C Ablation Analysis
We design ablation experiments and perform data analysis and visualization analysis on the results, respectively.
@@ V-C1 Data Analysis
The impact of incorporating OKD with mutual learning and AMB in multimodal learning is investigated through a detailed ablation analysis. Specifically, the role of each component is examined by progressively simplifying the proposed AMBOKD approach. MMOKD-DG and MMOKD-DK are the varients of our method, which keep the part of dynamic weights block and the dynamic gradients block in the AMB module, seprately. Then, MMOKD omits the entire AMB component, serving as a baseline to evaluate the underlying OKD and mutual learning mechanism. MKD further eliminates the online training method, representing a multi-teacher KD framework. The V-KD and E-KD approaches are further simplifications, using single-teacher KD from the visual and EEG modalities, respectively. Lastly, AMM is the most basic form of our model, which depends solely on a multi-head self-attention mechanism for modality fusion, without any KD. In addition, our analysis incorporates the unimodality encoders MCGRAM for EEG and EfficientNet for the visual domain.
The ablation results are presented in Table IV. As can be seen, these uni-modality encoders MCGRAM and EfficientNet are outperformed by the AMM model, which leverages multi-head self-attention for fusion, thereby validating the multi-head self-attention mechanism’s capacity to effectively integrate multimodal data. The results of E-KD, V-KD and M-KD further verify the effectiveness of the knowledge distillation mechanism to improve the performance of the multimodal fusion model. MMOKD outperform MKD by 1.37% and 1.9% in AUC, which demonstrate the effectiveness of mutual learning method for achieving fully modal interaction and improving multimodal model performance. The results of the MMOKD-DG and MMOKD-DK methods are further improved compared with the MMOKD method, which indicates that the adaptive balance of gradients and the interaction between different modal models can provide a more effective training scheme Our latest method, AMBOKD fuses the above two dynamic balancing schemes and achieves the optimal results, 2.08% and 2.05% higher than MMOKD, demonstrating the significant potential of modal balancing in multimodal mutual learning tasks.
@@
MCGRAM 82.47/7.11 $<10^{-3}$ 80.44/5.32 $<10^{-3}$
EfficientNet 78.81/8.89 $<10^{-3}$ 84.13/3.92 $<10^{-3}$
AMM 87.16/3.62 $<10^{-3}$ 87.75/3.64 $<10^{-3}$
EKD 88.48/4.28 $<10^{-3}$ 88.68/3.73 $<10^{-3}$
VKD 88.08/4.20 $<10^{-3}$ 87.53/4.76 $<10^{-3}$
MKD 90.21/4.45 $<10^{-3}$ 89.35/3.54 $<10^{-3}$
MMOKD 91.58/3.89 $<10^{-3}$ 91.25/2.47 $<10^{-3}$
MMOKD-DK 91.90/3.48 $<10^{-3}$ 91.61/2.92 $<10^{-3}$
MMOKD-DG 92.31/3.54 $<10^{-3}$ 92.59/2.21 $<10^{-3}$
AMBOKD (ours) 93.66/2.95 -- 93.30/2.64 --
Figure 11: The dynamic learning progress ratio changes of MMOKD, MMOKD-DG and AMBOKD in training process.

### v3

@@ V Experimental Results
In this section, the experimental settings with the parameter configurations of our proposed approach are first presented. Subsequently, a comprehensive comparison and ablation study is conducted to evaluate the performance of the proposed AMBOKD approach in the dim object recognition task. System verification through simulated real-world scenarios is designed to further validate the effectiveness of the system for dim object detection.
@@ V-A1 Configurations
The parameters of the visual and the EEG encoders are set to follow the configurations as described in the previous study [18] [5]. The number of heads is set to 2 in the multi-head self-attention module of the fusion model. The kernel sizes of the FC layers for alignment are configured as $1280\times 64$ and $256\times 64$ for visual and EEG features, respectively. Before computing the Q, K, and V matrices, the matrices in the FC layer have a size of $64\times 64$. The FC layers have a size of $128\times 2$ in the classifier of the Fusion domain.
The temperature coefficient $\tau$ is set to 4 in the OKD approaches to control the softness of the KD targets. The hyperparameters $\alpha$ and $\beta$ are dynamically adjusted by the AMB module, which contributes to the regularization component of the overall loss function. The hyperparameter $\gamma$ is set to 3 to fine-tune the sensitivity of the AMB module. In addition, we impose lower and upper limits, $R_{\min}$ and $R_{\max}$ at 0.1 and 10, respectively, to confine the range of dynamic modulation ratio $R^{DG}_{(t)}$ during the training process.
In the experiments, all models are implemented using the PyTorch framework. The Adam optimizer is employed for parameter optimization with a learning rate of $10^{-4}$. The number of training epochs is set to 15, and the batch size is 64. The optimization loss function consists of cross-entropy (CE) loss and knowledge distillation (KD) loss, with specific configurations varying by model.
To evaluate model performance, we adopt a cross-validation approach. Specifically, for each validation fold, the data from one subject is used as the validation set, while the remaining nine subjects’ data serve as the training set. This process is repeated across all subjects, and each fold is further evaluated using five different random seeds. The final results are obtained by averaging the performance across all folds and seeds.
@@ V-A2 Performance Metrics
A comprehensive set of metrics, including area under the receiver operating characteristic curve (AUC), accuracy (ACC), F1 score (F1) and Precision is used to evaluate the performance of classification models. Each metric is analyzed in terms of mean and standard deviation, offering insight into the reliability of the model and its ability to generalize across various datasets or conditions. Specifically, the AUC is emphasized as the main metric because of its effectiveness in measuring the discriminative ability of a model, especially in scenarios with unbalanced class distributions. To evaluate the object detection performance of the brain-eye-computer based system, AP@50:95 and AR@50:95 are used to measure the average precision and recall over Intersection Over Union (IOU) thresholds ranging from 0.50 to 0.95, while AP@50 and AR@50 are used for an IOU of 0.50. In addition, F1@50:95 and F1@50 provide a balanced view of accuracy by considering both precision and recall at these thresholds.
Additionally, to determine the statistical significance of performance differences, we first conduct a one-way analysis of variance (ANOVA) to assess significant differences among all models. Since we focus specifically on comparing our proposed model against other baseline models, we further conduct pairwise t-tests between the proposed model and each baseline model. The significance threshold is set at $p<0.05$. To address potential multiple comparison issues, Bonferroni correction is applied to adjust the p-values and control the family-wise error rate.
@@ V-B1 Comparison Results on ESSVP Dataset
To verify the effectiveness of the proposed AMBOKD approach, we conduct subject-independent comparison experiments against baseline and state-of-the-art models. Single-modal EEG methods include EEGNet [43] and MCGRAM [5], while single-modal vision methods include ResNet-50 [16] and EfficientNet [18]. A brief overview of the compared multimodal methods is as follows:
  • MKD [33]: A multimodal knowledge distillation approach that transfers knowledge from pre-trained EEG and visual teacher models.
  • EMKD [34]: An entropy-based adaptive multi-teacher knowledge distillation method.
  • CA-MKD [15]: A multi-teacher knowledge distillation technique that adaptively weights teacher predictions based on sample-wise confidence.
  • C2KD [44]: A cross-modal knowledge distillation method that uses bidirectional distillation, on-the-fly selection to filter misaligned samples, and proxy models to transfer knowledge across modalities.
  • DIST+ [45]: A knowledge distillation method using correlation-based loss and semantic similarities, with teacher acclimation and feature-level loss extension.
  • DML [32]: An online knowledge distillation method where multiple student models learn from each other by aligning predictive distributions.
  • KDCL [46]: An online knowledge distillation method that integrates contrastive learning to improve feature representation.
  • EML [35]: A method combining mutual learning and ensemble learning, where student models distill knowledge from both individual peers and an ensemble of predictions.
  • MLEMKD [47]: A mutual learning-empowered knowledge distillation framework that uses a soft-label evaluation mechanism to mitigate redundancy and improve knowledge transfer.
In addition, the EEG and visual modal model trained in AMBOKD are named AMBOKD-E and AMBOKD-V.
As indicated in Table I, our proposed AMBOKD achieves impressive results in the dim object recognition task of subject-independent experiments. This outcome surpasses the performance of baseline and state-of-the-art approaches. Additionally, the mutual learning mode with the AMB module enhances the performance of EEG and visual modality models: AMBOKD-E and AMBOKD-V outperform the original unimodal models MCGRAM and EfficientNet across AUC, ACC, and F1 metrics. The AUC-based ANOVA statistical test findings demonstrate that our proposed AMBOKD, AMBOKD-E, and AMBOKD-V significantly outperform all comparison approaches in subject-independent experiments ($p<0.05$).
@@ V-B2 Comparison Results on CIFAR-100 Dataset
We use the same training setup as described in [49], employing an SGD optimizer with a momentum of 0.9, a batch size of 64, a weight decay of $5\times 10^{-4}$. The initial learning rate is set at 0.05, with a decay factor of 10 every 30 epochs starting from epoch 150. For the AMB module, we specify a learning rate of $1\times 10^{-4}$. The comparison results with three different random seeds, presented as acc(@1, mean/standard deviation), are shown in Table II. Our AMBOKD achieve the best performance among all compared methods, demonstrating its robustness and practicability.
TABLE II: Comparison results on CIFAR-100 Test set
@@
V-B3 Comparison Results on SEED-VIG Dataset
To evaluate the effectiveness and generalizability of our proposed method, we conducted experiments using the publicly available multimodal SEED-VIG dataset [53]. SEED-VIG is a comprehensive dataset designed for vigilance estimation in simulated driving scenarios. The dataset quantifies drivers’ psychological fatigue states using the PERCLOS annotation method, which ranges from 0 to 1, with a threshold of 0.35 used to distinguish between awake and fatigued states for binary classification. It includes 21 channels of EEG data and 4 channels of EOG data collected from 23 subjects, with each subject contributing 885 samples.
[…]
The learning rate was set to 0.001, and the Adam optimizer with default parameter values was employed. The training process consisted of 30 epochs with a batch size of 64. A leave-one-out cross-validation approach was utilized for cross-subject experiments to evaluate the model’s generalization performance. EEGNet [43] and EOGNet [54] were used as unimodal models and as feature extractors for the multimodal model. The comparative results, averaged over five different random seeds, are presented in Table III. The results demonstrate that multimodal methods further improve performance metrics, with our proposed AMBOKD method achieving state-of-the-art performance, highlighting its effectiveness and generalizability in multimodal learning.
TABLE III: Comparison results on SEED-VIG dataset
Method AUC/Std(%) F1/Std(%) ACC/Std(%) Precision/Std(%)
EEGNet 83.48/0.60 75.29/0.36 73.76/0.48 82.60/0.44
EOGNet 89.96/0.09 80.13/0.76 78.43/0.79 86.32/0.44
MKD 91.31/0.41 79.67/0.47 78.47/0.66 86.60/0.23
EMKD 91.36/0.44 79.70/0.46 78.51/0.64 86.62/0.21
CA-MKD 91.31/0.45 79.59/0.47 78.39/0.65 86.58/0.20
C2KD 90.94/0.18 81.88/0.24 80.37/0.30 86.95/0.17
DIST+ 91.26/0.35 83.36/0.73 81.78/0.82 88.08/0.33
DML* 90.15/0.26 81.37/0.57 79.96/0.61 86.55/0.23
KDCL* 89.63/0.39 81.16/0.56 79.66/0.62 85.96/0.43
EML* 90.86/0.30 81.59/0.29 80.05/0.27 87.01/0.24
MLEMKD* 90.73/0.37 81.85/0.28 80.35/0.36 86.98/0.29
AMBOKD (ours)* 91.64/0.16 84.34/0.53 83.32/0.63 88.04/0.24
Figure 9: Dynamic optimization degree changes of MMOKD, MMOKD-DG and AMBOKD during training process on the ESSVP dataset.
@@ V-C Ablation Analysis
We conduct ablation experiments, performing data and visualization analyses on the results.
@@ V-C1 Data Analysis
The impact of incorporating OKD with mutual learning and AMB in multimodal learning is investigated through a detailed ablation analysis. Specifically, the role of each component is examined by progressively simplifying the proposed AMBOKD approach. MMOKD-DG and MMOKD-DK are the variants of our method, which keep the part of the dynamic weights block and the dynamic gradients block in the AMB module, separately. Then, MMOKD omits the entire AMB component, serving as a baseline to evaluate the underlying OKD and mutual learning mechanism. MKD further eliminates the online training method, representing a multi-teacher KD framework. The V-KD and E-KD approaches are further simplifications, using single-teacher KD from the visual and EEG modalities, respectively. Lastly, AMM is the most basic form of our model, which depends solely on a multi-head self-attention mechanism for modality fusion, without any KD. In addition, our analysis incorporates the unimodality encoders MCGRAM for EEG and EfficientNet for the visual domain.
The ablation results are presented in Table IV. As can be seen, these uni-modality encoders MCGRAM and EfficientNet are outperformed by the AMM model, which leverages multi-head self-attention for fusion, thereby validating the multi-head self-attention mechanism’s capacity to effectively integrate multimodal data. The results of E-KD, V-KD, and M-KD further verify the effectiveness of the knowledge distillation mechanism to improve the performance of the multimodal fusion model. MMOKD outperforms MKD by 1.37% and 1.9% in AUC, which demonstrates the effectiveness of the mutual learning method for achieving fully modal interaction and improving multimodal model performance. The MMOKD-DG and MMOKD-DK methods outperform the MMOKD method, indicating that adaptive gradient balancing and interactions between modality models provide a more effective training scheme. Our latest method, AMBOKD fuses the above two dynamic balancing schemes and achieves the optimal results, 2.08% and 2.05% higher than MMOKD, demonstrating the significant potential of modal balancing in multimodal mutual learning tasks.
@@
MCGRAM 82.47/7.11 $10^{-3}$ 80.44/5.32 $10^{-3}$
EfficientNet 78.81/8.89 $10^{-3}$ 84.13/3.92 $10^{-3}$
AMM 87.16/3.62 $10^{-3}$ 87.75/3.64 $10^{-3}$
EKD 88.48/4.28 $10^{-3}$ 88.68/3.73 $10^{-3}$
VKD 88.08/4.20 $10^{-3}$ 87.53/4.76 $10^{-3}$
MKD 90.21/4.45 $10^{-3}$ 89.35/3.54 $10^{-3}$
MMOKD 91.58/3.89 $10^{-3}$ 91.25/2.47 $10^{-3}$
MMOKD-DK 91.90/3.48 $10^{-3}$ 91.61/2.92 $10^{-3}$
MMOKD-DG 92.31/3.54 $10^{-3}$ 92.59/2.21 $10^{-3}$
AMBOKD (ours) 93.66/2.95 - 93.30/2.64 -